A well-known piano manufacturer wishes to expand in China. It decides that 80% of the $32 million it needs will come from debt, and the remaining 20% from selling equity. The cost of debt is 7% and the corporate tax is 35%. To estimate the cost of equity, the firm uses the CAPM with these parameters:

equation :

qi is the expected return on equity oee is the yield on a 3 month Treasury bill β is a coefficient of volatility of the companys returns relative to the market; a value of 1 is average (see week 3) ranker is the average market return in the sector the company operates and (rnarei- Tir) is the risk premium of that sector ,equity-rfree + β *(r,market-rfree), where where r,re=0.02, β=1.20, and-ket-012

Specifically I need help with these two elements:

1.What is the WACC, or hurdle rate, of the investment?

2. If the IRR of the investment is estimated at 7 %, should it proceed?
Answer #1

Computation of WACC or hurdle rate:

Step: 1 Computation of cost of Equity:

Cost of Equity = Risk-Free Rate + Beta X (Market Rate of Return - Risk-Free Rate)

                                = 0.02+1.2 X (0.12-0.02)

                                =0.02+0.12

Cost of Equity    =0.14

Cost of Debt       = 7% or 0.07 (Given in problem)

Step: 2 Computation of debt and equity amount:

Total cost of the project = 32 million

  • Proceeds from Equity = 32 million X 20 % =6.4 million
  • Proceeds from Debt    = 32 million X 80% =25.6 million.

Formula for WACC

WACC = ((E/V) X Re) + [((D/V) X Rd) X (1-T)]

Where

E= Market value of the company’s equity

D= Market value of the company’s debt

V= E+D

Re= Cost of Equity

Rd= Cost of Debt

T=tax rate

WACC = ((6.4 /32) X0.14) + [((25.6 /32) X0.07) X (1-35%)]

= 0.028+0.0364

= 0.0644

= 6.44%

  1. WACC or hurdle rate is = 6.44%
  1. If the IRR of the investment is estimated at 7% and the WACC is 6.44% so, it’s better to proceed.